# Geometric integrators for adiabatically closed simple thermodynamic systems

### v1

@@ 2.2 The continuous model: adiabatically closed simple thermodynamic systems
The vector field $\xi_{L}$ above is an example of SODE (Second Order Differential Equation), sometimes called a semispray or semigerbe. A SODE $\xi$ is a vector field on ${\mathsf{T}}Q$ such that its integral curves are just the tangent lifts of its projections to ${\mathsf{T}}Q$. Thus, in bundle coordinates $(q^{i},v^{i})$ on ${\mathsf{T}}Q$, a SODE reads
@@ 3.1 Discrete variational principle
In this section, we will construct a discrete model inspired in the continuous one presented in the precedent section.
@@
A discrete thermodynamic path, $\sigma\in\Omega_{T}\left(\{t_{k}\}_{k=0}^{N}\right)$, is said to be a solution of the discrete thermodynamic Euler–Lagrange equations if for all variations $\delta\sigma\in\mathcal{D}_{\sigma}$ with fixed extreme points, i.e., $\delta q_{0}=0=\delta q_{N}$, we have
@@ 3.2 Discrete flows
Let $L_{d}:Q\times Q\times\mathbb{R}\longrightarrow\mathbb{R}$ be a discrete Lagrangian such that ${\mathsf{D}}_{S}L_{d}$ does not vanish. The discrete thermodynamic Legendre transforms are the maps $\mathbb{F}^{f\pm}L_{d}:Q\times Q\times\mathbb{R}\longrightarrow{\mathsf{T}}^{\ast}Q\times\mathbb{R}$ given by
@@
$=\Bigg(q_{1},{\mathsf{D}}_{2}L_{d}\left(q_{0},q_{1},S_{0}\right)+\frac{1}{2}f^{\mathrm{fr,+}}_{d}\left(q_{0},q_{1},S_{0}\right),S_{0}+\frac{f^{\mathrm{fr,+}}_{d}\left(q_{0},q_{1},S_{0}\right)q_{1}-f^{\mathrm{fr,-}}_{d}\left(q_{0},q_{1},S_{0}\right)q_{0}}{{\mathsf{D}}_{S}L_{d}\left(q_{0},q_{1},S_{0}\right)}\Bigg),$
$\mathbb{F}^{f-}L_{d}\left(q_{0},q_{1},S_{0}\right)=\left(q_{0},-{\mathsf{D}}_{1}L_{d}\left(q_{0},q_{1},S_{0}\right)-\frac{1}{2}f^{\mathrm{fr,-}}_{d}\left(q_{0},q_{1},S_{0}\right),S_{0}\right).$
The second components of these maps, multiplied by the time step $h$, will be called the discrete momenta, namely
$p^{+}_{d}(q_{0},q_{1},S_{0})=h{\mathsf{D}}_{2}L_{d}\left(q_{0},q_{1},S_{0}\right)+\frac{h}{2}f^{\mathrm{fr,+}}_{d}\left(q_{0},q_{1},S_{0}\right),$ (22)
$p^{-}_{d}(q_{0},q_{1},S_{0})=-h{\mathsf{D}}_{1}L_{d}\left(q_{0},q_{1},S_{0}\right)-\frac{h}{2}f^{\mathrm{fr,-}}_{d}\left(q_{0},q_{1},S_{0}\right).$
@@
${\mathsf{D}}_{2}{\mathsf{D}}_{1}L_{d}-\frac{1}{2}{\mathsf{D}}_{2}f^{\mathrm{fr,-}}_{d},$
@@
$\left({\mathsf{D}}_{1}{\mathsf{D}}_{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{\mathrm{fr,+}}_{d}\right)\left(\left({\mathsf{D}}_{S}L_{d}\right)^{2}+\left({\mathsf{D}}_{S}f^{\mathrm{fr,+}}_{d}q_{2}-{\mathsf{D}}_{S}f^{\mathrm{fr,-}}_{d}q_{1}\right){\mathsf{D}}_{S}L_{D}-\left(f^{\mathrm{fr,+}}_{d}q_{2}-f^{\mathrm{fr,-}}_{d}q_{1}\right){\mathsf{D}}_{S}^{2}L_{d}\right)$-
$-\left({\mathsf{D}}_{2}{\mathsf{D}}_{S}L_{d}+\frac{1}{2}{\mathsf{D}}_{S}f^{\mathrm{fr,+}}_{d}\right)\left(\left(\left({\mathsf{D}}_{1}f^{\mathrm{fr,+}}_{d}\right)q_{2}-\left({\mathsf{D}}_{1}f^{\mathrm{fr,-}}_{d}\right)q_{1}-f^{\mathrm{fr,-}}_{d}\right){\mathsf{D}}_{S}L_{d}-\left(f^{\mathrm{fr,+}}_{d}q_{2}-f^{\mathrm{fr,-}}_{d}q_{1}\right){\mathsf{D}}_{1}{\mathsf{D}}_{S}L_{d}\right)\,.$
@@
Let $L_{d}:Q\times Q\times\mathbb{R}\longrightarrow\mathbb{R}$ be a discrete hyperregular Lagrangian. Then $\sigma=(q_{0},\ldots,q_{N},S_{0},S_{N})\in\Omega\left(\{t_{k}\}_{k=0}^{N}\right)$ is a solution of the discrete thermodynamic Euler–Lagrange equations if and only if
@@
If $\sigma\in\Omega\left(\{t_{k}\}_{k=0}^{N}\right)$ is a solution of the discrete thermodynamic Euler–Lagrange equations, we know that for every $k=1,\ldots,N-1$, we have $\mathbb{F}^{f-}L_{d}\left(q_{k},q_{k+1},S_{k}\right)=\mathbb{F}^{f+}L_{d}\left(q_{k-1},q_{k},S_{k-1}\right)$. Since $\mathbb{F}^{f-}L_{d}$ is a diffeomorphism, composing with its inverse, we get that $\Phi_{d}\left(q_{k-1},q_{k},S_{k-1}\right)=\left(q_{k},q_{k+1},S_{k}\right)$.
Conversely, if (23) holds, then composing to the left with $\mathbb{F}^{f-}L_{d}$ we get that the momentum matching equation holds and, thus, $\sigma$ is a solution of the discrete thermodynamic Euler–Lagrange equations. ∎
@@
$\left(\mathbb{F}^{f-}L_{d}\right)^{*}\mathrm{d}p=-\left({\mathsf{D}}_{1}^{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{\mathrm{fr,-}}_{d}\right)\mathrm{d}q_{1}-\left({\mathsf{D}}_{2}{\mathsf{D}}_{1}L_{d}+\frac{1}{2}{\mathsf{D}}_{2}f^{\mathrm{fr,-}}_{d}\mathrm{d}q_{2}\right)\,$,
$\left(\mathbb{F}^{f+}L_{d}\right)^{*}\mathrm{d}p=\left({\mathsf{D}}_{1}{\mathsf{D}}_{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{\mathrm{fr,+}}_{d}\right)\mathrm{d}q_{1}-\left({\mathsf{D}}_{2}^{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{2}f^{\mathrm{fr,+}}_{d}\mathrm{d}q_{2}\right)\,$,
@@
$\omega^{+}=\left({\mathsf{D}}_{1}{\mathsf{D}}_{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{\mathrm{fr,+}}_{d}\right)\mathrm{d}q_{1}\wedge\mathrm{d}q_{2}\,,\qquad\omega^{-}=\left({\mathsf{D}}_{1}{\mathsf{D}}_{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{\mathrm{fr,-}}_{d}\right)\mathrm{d}q_{1}\wedge\mathrm{d}q_{2}\,.$
@@
Given a hyperregular Lagrangian $L_{d}:Q\times Q\times\mathbb{R}\longrightarrow\mathbb{R}$, we have
$\Phi_{d}^{*}\,\omega^{-}=\omega^{+},\qquad\Phi_{d}^{*}\,\omega^{+}=\omega^{-}.$
@@
For the second equation, fixing $S_{0}\in\mathbb{R}$, we define the restricted discrete action map as $\mathcal{A}_{d}^{r}:Q\times Q\longrightarrow\mathbb{R}$ given by
@@ 3.3 Discrete Symmetries and Discrete Noether’s theorem
    $\xi_{Q\times Q\times\mathbb{R}}(L_{d})+f^{\mathrm{fr}}_{d}(\xi_{Q\times Q\times\mathbb{R}})=0$.
@@ 4.1 Damped harmonic oscillator
The corresponding thermodynamic Euler–Lagrange equations (10) read
@@
The discrete thermodynamic Euler–Lagrange equations (21) for this system provide the following integration method:
@@ 4.2 Ideal gas in a piston
We also consider a friction force between the cylinder and the piston, which is proportional to its velocity. In this case, we consider a constant temperature independent coefficient of friction $\gamma$:
@@
and the friction forces by
@@ 4.3 Van der Waals gas in a piston
In this setting, the thermodynamic Euler–Lagrange equations are the following:
@@
and the friction forces as follows:

### v2

@@ 1 Introduction
It is also worth mentioning the numerical methods for a supercritical van der Waals fluid enclosed in a cylinder developed by Fülöp, Szücz and Takács in [F.S.T2025]. Their approach is based on the GENERIC framework, which essentially consists of expressing a dissipative dynamical system in the form
$\frac{\mathrm{d}x^{i}}{\mathrm{d}t}=\sum_{j}\left(L^{i}_{j}\frac{\partial E}{\partial x^{j}}+M^{i}_{j}\frac{\partial S}{\partial x^{j}}\right)\,$, (1)
where $(L^{i}_{j})$ is a skew-symmetric matrix (representing a presymplectic form), $(M^{i}_{j})$ is a symmetric matrix, while $E$ and $S$ are functions that can be regarded as the total energy and the total entropy of the system, respectively. The authors then consider a generalisation of the symplectic Euler scheme, to discretise the evolution equations of the fluid. They obtain meritorious numerical results. However, they do not develop a general method for discretising thermodynamic systems, but rather construct an ad hoc generalisation of the symplectic method that serves their system under consideration. Moreover, their approach does not consider the geometry of the thermodynamic phase space whatsoever. We would also like to mention that the dynamics (1) corresponds to a Rayleigh-type dissipation, i.e. a dissipation proportional to velocity or momentum. Indeed, the symmetric matrix $(M^{i}_{j})$ determines a Rayleigh tensor. We previously studied systems with Rayleigh dissipation in [d.L.L2021, d.L.L2022] and their discretisation in [d.L.L2022a].
@@ 2.2 The continuous model: adiabatically closed simple thermodynamic systems
The vector field $\xi_{L}$ above is an example of SODE, sometimes called a semispray or semigerbe. A SODE $\xi$ is a vector field on ${\mathsf{T}}Q$ such that its integral curves are just the tangent lifts of its projections to ${\mathsf{T}}Q$. Thus, in bundle coordinates $(q^{i},v^{i})$ on ${\mathsf{T}}Q$, a SODE reads
@@ 3.1 Discrete variational principle
In this section, we will construct a discrete model inspired by the continuous one presented in the previous section. For simplicity’s sake, we will first consider the case without external forces. Discrete external forces will be added at the end of the present subsection.
@@
A discrete thermodynamic path $\sigma\in\Omega_{T}\left(\{t_{k}\}_{k=0}^{N}\right)$ is said to be a solution of the discrete thermodynamic Euler–Lagrange equations if for all variations $\delta\sigma\in\mathcal{D}_{\sigma}$ with fixed extreme points, i.e., $\delta q_{0}=0=\delta q_{N}$, we have
@@
We will now incorporate discrete external forces to the constructions above. Let $\left(Q,L_{d},f^{\mathrm{fr}}_{d}\right)$ be a simple discrete thermodynamic system. The discrete version of external forces is represented by a $1$-form $f^{\mathrm{ext}}_{d}=\left(f^{\mathrm{ext,-}}_{d},f^{\mathrm{ext,+}}_{d},0\right)\in\Omega^{1}(Q\times Q\times\mathbb{R})$. The $4$-tuple $\left(Q,L_{d},f^{\mathrm{fr}}_{d},f^{\mathrm{ext}}_{d}\right)$ will be called a simple discrete forced thermodynamic system.
Definition 6.
A discrete thermodynamic path $\sigma\in\Omega_{T}\left(\{t_{k}\}_{k=0}^{N}\right)$ is called a solution of the discrete thermodynamic forced Euler–Lagrange equations if for all variations $\delta\sigma\in\mathcal{D}_{\sigma}$ with fixed extreme points, i.e., $\delta q_{0}=0=\delta q_{N}$, we have
$\mathrm{d}\mathcal{A}_{d}\cdot\delta\sigma+\frac{1}{2}\sum_{k=1}^{N}\Big(f^{\mathrm{ext,-}}_{d}\left(q_{k},q_{k+1},S_{k}\right)\cdot\delta q_{k}+f^{\mathrm{ext,+}}_{d}\left(q_{k},q_{k+1},S_{k}\right)\cdot\delta q_{k+1}\Big)=0\,.$
To simplify the expressions, we introduce the notation
[…]
Using Theorem 6, we can obtain the following characterization of the solutions of the discrete thermodynamic forced Euler–Lagrange equations.
Corollary 8.
A discrete path $\sigma\in\Omega\left(\{t_{k}\}_{k=0}^{N}\right)$ is a solution of the discrete thermodynamic forced Euler–Lagrange equations if an only if it satisfies
${\mathsf{D}}_{1}L_{d}\left(q_{k},q_{k+1},S_{k}\right)+\frac{1}{2}f^{-}_{d}\left(q_{k},q_{k+1},S_{k}\right)+{\mathsf{D}}_{2}L_{d}\left(q_{k-1},q_{k},S_{k-1}\right)+\frac{1}{2}f^{+}_{d}\left(q_{k-1},q_{k},S_{k-1}\right)=0,$ (23a)
and
for every $k=1,\ldots,N$.
We shall refer to (23) as the discrete thermodynamic forced Euler–Lagrange equations.
@@ 3.2 Discrete flows
Let $\left(Q,L_{d},f^{\mathrm{fr}}_{d},f^{\mathrm{ext}}_{d}\right)$ be a simple discrete forced thermodynamic system. Assume that ${\mathsf{D}}_{S}L_{d}$ does not vanish. The discrete thermodynamic Legendre transforms are the maps $\mathbb{F}^{f\pm}L_{d}:Q\times Q\times\mathbb{R}\longrightarrow{\mathsf{T}}^{\ast}Q\times\mathbb{R}$ given by
@@
$=\Bigg(q_{1},{\mathsf{D}}_{2}L_{d}\left(q_{0},q_{1},S_{0}\right)+\frac{1}{2}f^{+}_{d}\left(q_{0},q_{1},S_{0}\right),S_{0}+\frac{f^{\mathrm{fr,+}}_{d}\left(q_{0},q_{1},S_{0}\right)q_{1}-f^{\mathrm{fr,-}}_{d}\left(q_{0},q_{1},S_{0}\right)q_{0}}{{\mathsf{D}}_{S}L_{d}\left(q_{0},q_{1},S_{0}\right)}\Bigg),$
$\mathbb{F}^{f-}L_{d}\left(q_{0},q_{1},S_{0}\right)=\left(q_{0},-{\mathsf{D}}_{1}L_{d}\left(q_{0},q_{1},S_{0}\right)-\frac{1}{2}f^{-}_{d}\left(q_{0},q_{1},S_{0}\right),S_{0}\right)\,$,
where $f^{\pm}_{d}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=f^{\mathrm{fr,\pm}}_{d}+f^{\mathrm{ext,\pm}}_{d}$. The second components of these maps, multiplied by the time step $h$, will be called the discrete momenta, namely
$p^{+}_{d}(q_{0},q_{1},S_{0})=h{\mathsf{D}}_{2}L_{d}\left(q_{0},q_{1},S_{0}\right)+\frac{h}{2}f^{+}_{d}\left(q_{0},q_{1},S_{0}\right),$ (24)
$p^{-}_{d}(q_{0},q_{1},S_{0})=-h{\mathsf{D}}_{1}L_{d}\left(q_{0},q_{1},S_{0}\right)-\frac{h}{2}f^{-}_{d}\left(q_{0},q_{1},S_{0}\right).$
@@
${\mathsf{D}}_{2}{\mathsf{D}}_{1}L_{d}+\frac{1}{2}{\mathsf{D}}_{2}f^{-}_{d},$
@@
$\left({\mathsf{D}}_{1}{\mathsf{D}}_{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{+}_{d}\right)\left(\left({\mathsf{D}}_{S}L_{d}\right)^{2}+\left({\mathsf{D}}_{S}f^{\mathrm{fr,+}}_{d}q_{2}-{\mathsf{D}}_{S}f^{\mathrm{fr,-}}_{d}q_{1}\right){\mathsf{D}}_{S}L_{D}-\left(f^{\mathrm{fr,+}}_{d}q_{2}-f^{\mathrm{fr,-}}_{d}q_{1}\right){\mathsf{D}}_{S}^{2}L_{d}\right)$-
$-\left({\mathsf{D}}_{2}{\mathsf{D}}_{S}L_{d}+\frac{1}{2}{\mathsf{D}}_{S}f^{+}_{d}\right)\left(\left(\left({\mathsf{D}}_{1}f^{\mathrm{fr,+}}_{d}\right)q_{2}-\left({\mathsf{D}}_{1}f^{\mathrm{fr,-}}_{d}\right)q_{1}-f^{\mathrm{fr,-}}_{d}\right){\mathsf{D}}_{S}L_{d}-\left(f^{\mathrm{fr,+}}_{d}q_{2}-f^{\mathrm{fr,-}}_{d}q_{1}\right){\mathsf{D}}_{1}{\mathsf{D}}_{S}L_{d}\right)\,.$
@@
Let $L_{d}:Q\times Q\times\mathbb{R}\longrightarrow\mathbb{R}$ be a discrete hyperregular Lagrangian. Then $\sigma=(q_{0},\ldots,q_{N},S_{0},S_{N})\in\Omega\left(\{t_{k}\}_{k=0}^{N}\right)$ is a solution of the discrete thermodynamic forced Euler–Lagrange equations if and only if
@@
If $\sigma\in\Omega\left(\{t_{k}\}_{k=0}^{N}\right)$ is a solution of the discrete thermodynamic forced Euler–Lagrange equations, we know that for every $k=1,\ldots,N-1$, we have $\mathbb{F}^{f-}L_{d}\left(q_{k},q_{k+1},S_{k}\right)=\mathbb{F}^{f+}L_{d}\left(q_{k-1},q_{k},S_{k-1}\right)$. Since $\mathbb{F}^{f-}L_{d}$ is a diffeomorphism, composing with its inverse, we get that $\Phi_{d}\left(q_{k-1},q_{k},S_{k-1}\right)=\left(q_{k},q_{k+1},S_{k}\right)$.
Conversely, if (25) holds, then composing to the left with $\mathbb{F}^{f-}L_{d}$ we get that the momentum matching equation holds and, thus, $\sigma$ is a solution of the discrete thermodynamic forced Euler–Lagrange equations. ∎
@@
$\left(\mathbb{F}^{f-}L_{d}\right)^{*}\mathrm{d}p=-\left({\mathsf{D}}_{1}^{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{-}_{d}\right)\mathrm{d}q_{1}-\left({\mathsf{D}}_{2}{\mathsf{D}}_{1}L_{d}+\frac{1}{2}{\mathsf{D}}_{2}f^{-}_{d}\right)\mathrm{d}q_{2}\,$,
$\left(\mathbb{F}^{f+}L_{d}\right)^{*}\mathrm{d}p=\left({\mathsf{D}}_{1}{\mathsf{D}}_{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{+}_{d}\right)\mathrm{d}q_{1}-\left({\mathsf{D}}_{2}^{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{2}f^{+}_{d}\right)\mathrm{d}q_{2}\,$,
@@
$\omega^{+}=\left({\mathsf{D}}_{1}{\mathsf{D}}_{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{+}_{d}\right)\mathrm{d}q_{1}\wedge\mathrm{d}q_{2}\,,\qquad\omega^{-}=\left({\mathsf{D}}_{1}{\mathsf{D}}_{2}L_{d}+\frac{1}{2}{\mathsf{D}}_{1}f^{-}_{d}\right)\mathrm{d}q_{1}\wedge\mathrm{d}q_{2}\,.$
@@
Let $\left(Q,L_{d},f^{\mathrm{fr}}_{d},f^{\mathrm{ext}}_{d}\right)$ be a simple discrete forced thermodynamic system. Assume that $L_{d}$ is hyperregular. Then,
$\Phi_{d}^{*}\,\omega^{-}=\omega^{+}\,.$
Moreover, if $f^{\mathrm{ext}}_{d}\equiv 0$, then
$\Phi_{d}^{*}\,\omega^{+}=\omega^{-}\,.$
@@
For the second equation, if $f^{\mathrm{ext}}_{d}=0$, fixing $S_{0}\in\mathbb{R}$, we define the restricted discrete action map as $\mathcal{A}_{d}^{r}:Q\times Q\longrightarrow\mathbb{R}$ given by
@@ 3.3 Discrete Symmetries and Discrete Noether’s theorem
    $\xi_{Q\times Q\times\mathbb{R}}(L_{d})+f_{d}(\xi_{Q\times Q\times\mathbb{R}})=0$.
@@ 4.1 Damped harmonic oscillator
We will consider no external forces act on the system. The corresponding thermodynamic Euler–Lagrange equations (11) read
@@
$f^{\mathrm{fr,-}}_{d}\left(q_{0},q_{1},S_{0}\right)=-\gamma\frac{q_{1}-q_{0}}{h}\mathrm{d}q_{0}\,,\quad f^{\mathrm{fr,+}}_{d}\left(q_{0},q_{1},S_{0}\right)=-\gamma\frac{q_{1}-q_{0}}{h}\mathrm{d}q_{1}\,$,
and take $f^{\mathrm{fr,\pm}}_{d}=0$. The discrete thermodynamic Euler–Lagrange equations (22) for this system provide the following integration method:
@@ 4.2 Ideal gas in a piston
We also consider a friction force between the cylinder and the piston, which is proportional to its velocity, and no external force acting on the piston. In this case, we consider a constant temperature independent coefficient of friction $\gamma$:
@@
the external forces $f^{\mathrm{ext,\pm}}_{d}=0$ and the friction forces by
@@ 4.3 Van der Waals gas in a piston
We will assume again no external forces act on the piston. In this setting, the thermodynamic Euler–Lagrange equations are the following:
@@
the external forces $f^{\mathrm{ext,\pm}}_{d}=0$ and the friction forces as follows:
@@
4.4 Expansion of a Van der Waals gas against external pressure
Most of the physically interesting examples of thermodynamic systems involve a non-zero external force. For the case of a Van der Waals gas contained in a piston, this is usually given by the external pressure.
To make a comparison of the results given by the discrete thermodynamic forced Euler-Lagrange equations and the ones obtained via the GENERIC-Motivated method introduced in [Fulop2025], we will work with the reduced version of the Van der Waals equations. The continuous Lagrangian for this system, considering a monoatomic gas with three degrees of freedom, is given by:
$L\left(V_{r},B_{r},S_{r}\right)=m_{p}\frac{B_{r}^{2}}{2A^{2}}-\left(\frac{2^{4}}{3V_{r}-1}\right)^{2/3}\mathrm{exp}\left(\left(S_{r}-1\right)\left(\frac{2}{3}\ln{2x_{c}}+\frac{5}{3}\right)\right)+\frac{3}{V_{r}},$
where $V_{r},B_{r},S_{r}$ are the reduced versions of the volume enclosed by the piston, the change rate of that volume in time and the entropy, $A$ is the area of the piston, $m_{p}$ its mass and $x_{c}$ is an adimensional coefficient that depends on the parameters of the gas being modeled, which is given by:
$x_{c}=\frac{T_{c}}{8p_{c}}\left(\frac{m_{\mathrm{gas}}T_{c}}{2\pi\hbar}\right)^{\frac{2}{3}}\,$,
where $m_{\mathrm{gas}}$ is the mass of a gas molecule. We have assumed that the gas in the piston is He-4 and used the values of its critical magnitudes and molar mass in [He41989]. From now on, we will consider, as in [Fulop2025], the quotient $\frac{m_{p}}{A^{2}}=1$.
$\tilde{F}^{\mathrm{fr}}=-\gamma B_{r}\mathrm{d}V_{r}\,.$
We will also consider an external force given by the reduced pressure of the exterior:
$\tilde{F}^{\mathrm{fr}}=-p_{A}\mathrm{d}V_{r}\,.$
$L_{d}\left(q_{0},q_{1},S_{0}\right)=\frac{\left(q_{1}-q_{0}\right)^{2}}{2h^{2}}-\left(\frac{2^{5}}{3\left(q_{1}+q_{0}\right)-2}\right)^{2/3}\mathrm{exp}\left(\left(S_{0}-1\right)\left(\frac{2}{3}\ln{2x_{c}}+\frac{5}{3}\right)\right)+\frac{6}{q_{1}+q_{0}}\,$,
and the friction and external forces as follows:
$f^{\mathrm{fr,-}}_{d}\left(q_{0},q_{1},S_{0}\right)=-\gamma\frac{q_{1}-q_{0}}{h}\,\mathrm{d}q_{0}\,$, $f^{\mathrm{fr,+}}_{d}\left(q_{0},q_{1},S_{0}\right)=-\gamma\frac{q_{1}-q_{0}}{h}\,\mathrm{d}q_{1}\,$,
$f^{\mathrm{ext,-}}_{d}\left(q_{0},q_{1},S_{0}\right)=-p_{A}\,\mathrm{d}q_{0}\,$, $f^{\mathrm{ext,+}}_{d}\left(q_{0},q_{1},S_{0}\right)=-p_{A}\,\mathrm{d}q_{1}\,.$
We have carried out the integration of the discrete thermodynamic forced Euler-Lagrange equations for this model using a time step $h=0.01$, taking $\gamma=0.1$, $p_{A}=1.8$ and for initial conditions $q_{0}=1.5$, $S_{0}=1.003702793854103$ and $B_{r}(0)=0$ and, thus $q_{1}=q_{0}$, which correspond to the initial conditions $V_{r}=1.5$, $T_{r}=1.01$.
The results have been compared with the ones given by the GENERIC-motivated extended symplectic method considered in [Fulop2025] (choosing the coefficients $\alpha_{b}=\alpha_{p}=1$ for such method), which was shown to outperform the classical Euler method.
In Figure 8 it can be seen that for the same time step, the integration of the discrete thermodynamic forced Euler-Lagrange equations outperforms the GENERIC-motivated algorithm, although a non linear equation must be solved at each step, which increases the computational demand.
A comparison when $p_{A}=0$ between these two methods is also shown in Figure 9.
(a) Reduced Volume
(b) Detailed section
Figure 8: Integration of $V_{r}$ for a Van der Waals gas composed of monoatomic molecules contained in a cylinder with a time step $h=0.01$ using the Discrete thermodynamic Euler–Lagrange equations solution (in green and dashed) and the GENERIC-motivated symplectic method (in blue and dotted), compared with the solution given by a RK-4 method with a tolerance of $10^{-15}$ (in black and solid).
(a) Reduced Volume
(b) Detailed section
Figure 9: Integration of $V_{r}$ for a Van der Waals gas composed of monoatomic molecules contained in a cylinder in absence of external forces with a time step $h=0.01$ using the Discrete thermodynamic Euler–Lagrange equations solution (in green and dashed) and the GENERIC-motivated symplectic method (in blue and dotted), compared with the solution given by a RK-4 method with a tolerance of $10^{-15}$ (in black and solid).